OpenStudy (anonymous):

find all real solutions: (1-x)^5-40=0

OpenStudy (anonymous):

you just need answer or you wanna know how to solve it..

OpenStudy (campbell_st):

well \[(1-x)^5 = 40\] \[(1 - x) = \sqrt[5]{40}\] you can finish it

OpenStudy (anonymous):

@ campbell_st then, you need show that it has only one real solution..